Grade 128 grade maths

Calculate the amount and compound interest on
₹ 62500 for 1whole 𝟏/ 𝟐 years at 8% per annum compounded half yearly.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Principal (P) = ₹ 62,500
Rate = 8% per annum or 4% per half year
Number of years =1 whole 1/2
There will be 3 half years in 1 whole 1/2year
Amount, A =P (1 +R/100 )^n
=₹[62500(1+4/100 )^3]
=₹ [62500 (104/100)^3]100
=₹ [62500 (26/25)^3]
=₹ 70304
C.I. = A − P = ₹ 70304 − ₹ 62500 = ₹ 7,804

Thanks